OpenStudy (magbak):

PLEASE HELP ME!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!! I WILL AWARD MEDAL!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!! What is the length of the major axis of the ellipse given by the equation? 8 9 16 18

OpenStudy (magbak):

\[(x^2/81)+(y^2/64)=1\]

OpenStudy (magbak):

@Loser66 @Hero @satellite73 @texaschic101

OpenStudy (magbak):

@satellite73 @satellite73 @satellite73 @satellite73

OpenStudy (anonymous):

hi

OpenStudy (magbak):

HI

OpenStudy (anonymous):

center of your ellipse is \((0,0)\) right?

OpenStudy (magbak):

if it is not in the question then it is not their.

OpenStudy (anonymous):

i am telling you it is \((0,0)\) because the general form is \[\frac{(x-h)^2}{a^2}+\frac{(y-k)^2}{b^2}=1\] and is this case \(h=k=0\) so the center is \((0,0)\)

OpenStudy (magbak):

OH ok.

OpenStudy (anonymous):

the major axis goes from \(-a\) to \(a\) here you have \(a^2=81\) so what is \(a\) ?

OpenStudy (magbak):

It is 9

OpenStudy (magbak):

is this the answer.

OpenStudy (anonymous):

hold on

OpenStudy (magbak):

yes

OpenStudy (anonymous):

it asks for the length of the major axis, which goes from \((-9,0)\) to \((9,0)\)

OpenStudy (magbak):

also I have ONLY ONE MORE QUESTION TO ASK YOU PLEASE PLEASE PLEASE PLEASE PLEASE PLEASE PLEASE PLEASE

OpenStudy (anonymous):

so the answer is NOT 9

OpenStudy (magbak):

OK so the answer can be found by using the distance formula?

OpenStudy (anonymous):

it is 18 the length of the "semi major" axis is 9

OpenStudy (anonymous):

you don't need the distance formula to find the distance between \((-9,0)\) and \((9,0)\)

OpenStudy (anonymous):



OpenStudy (anonymous):

lousy picture but you get the idea distance is \(18\)

OpenStudy (magbak):

yes of course.

OpenStudy (anonymous):

ok i can help you with the next one as well

OpenStudy (magbak):

ok thanks also I found one more question after that.